Solve the system of equations:
x + 2y = 3
4x + 5y = 6

Found 2 solutions by stanbon, msladyg32:
Answer by stanbon(75887)

Solve the system of equations:
x + 2y = 3
4x + 5y = 6
----------------------
Multiply thru the 1st eq. by 4 to get:
4x + 8y = 12
4x + 5y = 6
------
Subtract the 2nd from the 1st and solve for "y":
3y = 6
y = 2
----
Since x+ 2y = 3, x +2(2) = 3, x = -1
--------------------
Solution: (-1,2)
===================
Cheers,
Stan H.

Answer by msladyg32(1)

x + 2y = 3
4x + 5y = 6
x*5 + 2y*5 = 3*5
4x*2 + 5y * 2 = 6*2
5x + 10y = 15
8x + 10y = 12
-3x=3
-3x/-3 = 3/-3
x=-1